OpenStudy (anonymous):

Can I also get help with finding this integral : x^3/x dx

OpenStudy (anonymous):

\[ \Large \int \frac{x^3}{x}dx=\int x^2\ dx=\frac{x^3}{3} \]

OpenStudy (anonymous):

Or did you mean \[\int x^{3/x}dx\]?

OpenStudy (jit4won):

@nbouscal Can you solve the integral you wrote

OpenStudy (anonymous):

my bad I meant x^3 -7 / x dx

OpenStudy (jit4won):